OpenStudy (cenaida):

Use the distributive property to simplify the following expression: p(p2 + 3) + 2p A. p2 + 3p + 3 B. 2p2 + 2p + 3 C. p2 + 5p - 3 D. p3 + 5p

OpenStudy (cj49):

pxp^2=p^3 p^3+3p+2p=?

OpenStudy (cenaida):

A?

OpenStudy (anonymous):

\[\left(p^2+3\right) p+2 p=p^3+5 p \]D.

OpenStudy (cenaida):

thanks rob!